# Geodesic Slice Sampling on the Sphere

### v3

@@ Abstract
Probability measures on the sphere form an important class of statistical models and are used, for example, in modeling directional data or shapes. Due to their widespread use, but also as an algorithmic building block, efficient sampling of distributions on the sphere is highly desirable. We propose a shrinkage based and an idealized geodesic slice sampling Markov chain, designed to generate approximate samples from distributions on the sphere. In particular, the shrinkage-based version of the algorithm can be implemented such that it runs efficiently in any dimension and has no tuning parameters. We verify reversibility and prove that under weak regularity conditions geodesic slice sampling is uniformly ergodic. Numerical experiments show that the proposed slice samplers achieve excellent mixing on challenging targets including the Bingham distribution and mixtures of von Mises-Fisher distributions. In these settings our approach outperforms standard samplers such as random-walk Metropolis-Hastings and Hamiltonian Monte Carlo.
@@ 1 Introduction
In recent years, with the advent of sampling methods based on Markov chains, Bayesian inference with posterior distributions on manifolds attracted considerable attention. In particular, various Markov chain algorithms for approximate sampling on manifolds have been developed (see Byrne and Girolami, 2013; Lan et al., 2014; Goyal and Shetty, 2019; Lie et al., 2023; Beskos and Kamatani, 2022). Here we focus on the prototypical case of the sphere embedded in $\mathbb{R}^{d}$ with its inherent geometrical features as the underlying manifold. Following the slice sampling paradigm, we introduce and analyze an efficient way of approximate sampling of distributions on the sphere.
Let $\mathbb{S}^{d-1}$ be the $(d-1)$-dimensional Euclidean unit sphere in $\mathbb{R}^{d}$ and let $\sigma_{d-1}$ denote the volume measure on $\mathbb{S}^{d-1}$. For $p\colon\mathbb{S}^{d-1}\to(0,\infty)$ satisfying
@@
    Posteriors naturally defined on the sphere require efficient sampling: Sampling distributions on the sphere play an important role in directional statistics and shape analysis (see Mardia and Jupp, 2000), and Bayesian inverse problems on $\mathbb{S}^{2}$ occur, e.g., in astrophysics or geophysics (see Marignier et al., 2023). Moreover, Bayesian density estimation (Holbrook et al., 2020) requires sampling of a posterior on an infinite-dimensional sphere, which usually is approximated by truncating the dimension, ending up with $\mathbb{S}^{d-1}$ for large $d$.
@@
    Spherical sampling can be used as tool within transforming Markov chains: Lan et al. (2014) introduced a Hamiltonian Monte Carlo (HMC) scheme for sampling from spherical distributions that leads, after suitable transformations, to an efficient exploration of distributions in $\mathbb{R}^{d}$ constrained by inequalities. Recently, Yang et al. (2022) demonstrated that MCMC algorithms on $\mathbb{S}^{d-1}$ can target distributions in $\mathbb{R}^{d}$ by mapping them to the sphere by means of the stereographic projection. They report that for their proposed algorithms this approach is advantageous in heavy-tailed distribution scenarios.
We follow the slice sampling paradigm for approximately simulating $\pi$, since it allows us to avoid local random walk behavior, no tuning of a parameter such as a step-size is required and usually a richer choice of possible updates is offered, for details see Neal (2003); Murray et al. (2010). A transition mechanism for realizing a corresponding Markov chain works as follows: Given current state $x\in\mathbb{S}^{d-1}$, a superlevel set $L(t)$ of $p$ is randomly determined by choosing a level $t\in(0,p(x))$. Then, the next Markov chain instance is specified by (mimicking) sampling of the normalized reference measure $\sigma_{d-1}$ restricted to the level set. The latter step requires some care regarding its algorithmic design. The main contributions of our paper are:
  • We introduce an ideal geodesic slice sampler with corresponding Markov kernel $H$ for approximately simulating $\pi$. The kernel $H$ is implemented by moving along a randomly chosen great circle through the current state that forms a geodesic on $\mathbb{S}^{d-1}$. The resulting sampling on the geodesic intersected with the level set can be algorithmically realized by using an univariate acceptance/rejection approach.
@@
  • We test our algorithms on challenging targets such as the Bingham distribution and mixtures of von Mises-Fisher distributions. We observe that our slice samplers outperform standard samplers such as random walk Metropolis Hastings (RWMH) and Hamiltonian Monte Carlo (HMC) on spherical distributions.
Let us comment on how our contributions fit into the literature. Lie et al. (2023) present two MCMC algorithms to sample distributions on the sphere using push forward kernels of a preconditioned Crank-Nicolson algorithm and of elliptical slice sampling, respectively. The crucial difference to our work is that Lie et al. (2023) assume that the target density is defined relative to an angular Gaussian distribution. Instead, here we consider target densities relative to the volume measure of the sphere. Observe that with increasing dimension both reference measures become “increasingly singular” to each other. Furthermore, Marignier et al. (2023) propose a proximal MCMC method to sample from posterior distributions of inverse imaging problems on $\mathbb{S}^{2}$. An infinite-dimensional setting of Bayesian density estimation is treated by Holbrook et al. (2020) by an HMC algorithm on a sphere. Other MCMC approaches have been developed on more general manifolds that usually cover $\mathbb{S}^{d-1}$ as a special case. This includes the HMC algorithm for manifolds embedded in the Euclidean space introduced by Byrne and Girolami (2013) which uses the geodesic flow. Furthermore, Goyal and Shetty (2019) investigate a Metropolis-Hastings-like geodesic walk on manifolds with non-negative curvature. Here it is also worth to mention that several geodesic walks exist that target the uniform distribution on subsets of the sphere. We refer to Section 2.3 for more details. Finally, note that recently there has been some theoretical progress in the investigation of convergence properties of slice sampling (see for example Łatuszyński and Rudolf, 2024; Natarovskii et al., 2021b, a; Hasenpflug et al., 2024) that very much influenced the presentation and proof arguments of our theoretical results.
@@ 2 Preliminaries
In this section we provide some general notation, give a brief introduction to geometrical features and explain how these can be leveraged to “walk the sphere”.
@@ 2.1 Setting and Notation
[…]
@@
On generic Borel space $(\mathcal{X},\mathcal{B}(\mathcal{X}))$ define the restriction of a measure $\nu$ to a set $A\in\mathcal{B}(\mathcal{X})$ by $\nu|_{A}$, that is $\nu|_{A}(B)=\nu(A\cap B)$ for all $B\in\mathcal{B}(\mathcal{X})$. Whenever appropriate, for a set $A\in\mathcal{B}(\mathcal{X})$ we denote by $\mathcal{U}_{A}$ the uniform distribution on $A$. For example, for $\mathcal{X}=\mathbb{R}$ with $A\in\mathcal{B}(\mathbb{R})$ the probability measure $\mathcal{U}_{A}$ refers to $\frac{1}{\lambda(A)}\,\lambda|_{A}$ if $\lambda(A)>0$ and for $\mathcal{X}=\mathbb{S}^{d-1}$ the distribution $\mathcal{U}_{\mathbb{S}^{d-1}}$ refers to $\frac{1}{\omega_{d-1}}\sigma_{d-1}$, where $\omega_{d-1}\coloneqq\sigma_{d-1}\left(\mathbb{S}^{d-1}\right)$. Sometimes it is more convenient to work with random variables. Therefore, let $(\Omega,\mathcal{F},\mathbb{P})$ be a sufficiently rich probability space on which all random variables occurring in this paper are defined. If a random variable $X$ has distribution $\nu$, we write $X\sim\nu$.
@@ 2.2 The Geometry of the Sphere
the great subsphere with pole $x$. This $(d-2)$-dimensional subsphere is the intersection of $\mathbb{S}^{d-1}$ and the $(d-1)$-dimensional hyperplane perpendicular to $x$, and in this sense can be thought of as a generalization of the equator. We also equip ${\mathbb{S}^{d-2}_{x}}$ with the corresponding volume measure, which we denote as $\mu_{x}$. Since ${\mathbb{S}^{d-2}_{x}}$ is essentially “a tilted $\mathbb{S}^{d-2}$”, we have $\mu_{x}({\mathbb{S}^{d-2}_{x}})=\omega_{d-2}$ for all $x\in\mathbb{S}^{d-1}$, cf. Appendix A. Accordingly $\frac{1}{\omega_{d-2}}\mu_{x}$ coincides with $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}$. In our transition mechanism, sampling w.r.t. the latter distribution frequently appears, such that we provide a procedure for performing it in Appendix A.
@@
Intuitively, we can also think of $\gamma_{(x,v)}$ as the curve obtained by “moving from the point $x$ in the direction of $v$ with a constant velocity”. This interpretation originates from Riemannian geometry and following their terminology we use the names great circle and geodesic interchangeably for the object $\gamma_{(x,v)}$. Note that all great circles passing through a point $x\in\mathbb{S}^{d-1}$ are of the from $\gamma_{(x,v)}$ for some $v\in{\mathbb{S}^{d-2}_{x}}$, i.e., the geodesics through $x\in\mathbb{S}^{d-1}$ are parametrized by ${\mathbb{S}^{d-2}_{x}}$.
Observe, due to the periodicity of sine and cosine, that all great circles are $2\uppi$-periodic. This is exploited when we incorporate the geodesics into our slice sampling approach.
@@
for $\theta\in\mathbb{R}$, where for brevity we define $\mathcal{S}\mathbb{S}^{d-1}\coloneqq\bigcup_{x\in\mathbb{S}^{d-1}}\big{(}\{x\}\times{\mathbb{S}^{d-2}_{x}}\big{)}$. If we interpret for a pair $(x,v)\in\mathcal{S}\mathbb{S}^{d-1}$ the first component as position on the sphere and the second component as “direction of view”, then applying the map $T_{\theta}$ to $(x,v)$ can be thought of as “following the geodesic $\gamma_{(x,v)}$ from $x$ in direction $v$ for a length of $\theta$ and then doing a U-turn back on the spot”. This intuition is underpinned by the following lemma, proven in Appendix B, which essentially tells us that following the geodesic corresponding to the pair $T_{\theta}(x,v)$ is the same as following the original geodesic $\gamma_{(x,v)}$ in reverse direction with an offset of $\theta$.
@@
For the first step in the outlined transition we can use the fact that geodesics through $x\in\mathbb{S}^{d-1}$ are parametrized by ${\mathbb{S}^{d-2}_{x}}$. Namely, we realize “a random geodesic” as $\gamma_{(x,v)}$, where $v$ is a sample from $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}$, see Algorithm 5 in Appendix A for the simulation of this probability measure. The distribution of the randomly chosen point on the geodesic in step 2 significantly influences the stationary distribution of the random walk on the sphere. In this section, we consider it to be independent of $x$, which causes the stationary distribution to be the uniform one. In Section 3, we explain how the second step can be modified so as to obtain a Markov chain with a desired stationary probability measure.
Using a distribution $\tau$ on $[0,2\uppi)$ to sample on the geodesic, we obtain the following Markov kernel
@@
1:  Draw $V\sim\mathcal{U}_{\mathbb{S}^{d-2}_{x}}$, call the result $v$.   {Perform Algorithm 5 with input $x$.}
@@
For targeting the uniform distribution on the sphere, the geodesic random walk introduced by Goyal and Shetty (2019, Algorithm 1, for their $K=\mathbb{S}^{d-1}$) corresponds to $K_{\tau}$ with $\tau$ being the distribution of $\varepsilon R$, where $R$ is a chi-distributed random variable with $d-1$ degrees of freedom and fixed $\varepsilon>0$. For manifolds with non-negative sectional curvature and bounded Riemannian curvature tensor, Goyal and Shetty (2019) provide mixing time results if their unfiltered walk is modified to target the uniform distribution of a strongly convex subset of the ambient manifold.
@@ 3 Geodesic Slice Sampling in $\mathbb{S}^{d-1}$
Proof  By the lower semicontinuity $p^{-1}((t,\infty))=L(t)$ is open, and by the fact that $x\in L(t)$ it is also non-empty. Moreover, since $\gamma_{(x,v)}$ is continuous we have that $L(x,v,t)=\gamma_{(x,v)}^{-1}(L(t))$ is again open and non-empty, such that $\lambda(L(x,v,t))>0$. The final assertion follows by observing that $L(\left\|p\right\|_{\infty})$ is open (by the lower semicontinuity) and by the definition of $\left\|p\right\|_{\infty}$ satisfies $\sigma_{d-1}(L(\left\|p\right\|_{\infty}))=0$. Consequently, we have $L(\left\|p\right\|_{\infty})=\emptyset$, such that $\left\|p\right\|_{\infty}=\sup_{x\in\mathbb{S}^{d-1}}p(x)$.

### v4

@@ Abstract
Probability measures on the sphere form an important class of statistical models and are used, for example, in modeling directional data or shapes. Due to their widespread use, but also as an algorithmic building block, efficient sampling of distributions on the sphere is highly desirable. We propose a shrinkage based and an idealized geodesic slice sampling Markov chain, designed to generate approximate samples from distributions on the sphere. In particular, the shrinkage-based version of the algorithm can be implemented such that it runs efficiently and has no tuning parameters. We verify reversibility and prove that under weak regularity conditions geodesic slice sampling is uniformly ergodic. Numerical experiments show that the proposed slice samplers achieve excellent mixing on challenging targets including distributions arising in rigid-registration problems and mixtures of von Mises-Fisher distributions. In these settings our approach outperforms standard samplers such as random-walk Metropolis-Hastings and Hamiltonian Monte Carlo.
@@ 1 Introduction
In recent years, with the advent of sampling methods based on Markov chains, Bayesian inference with posterior distributions on manifolds attracted considerable attention. In particular, various Markov chain algorithms for approximate sampling on different manifolds have been developed (see Byrne and Girolami, 2013; Lan et al., 2014; Goyal and Shetty, 2019; Lie et al., 2023; Beskos and Kamatani, 2022). Here we focus on the prototypical case of the sphere embedded in $\mathbb{R}^{d}$ with its inherent geometrical features as the underlying manifold. Following the slice sampling paradigm, we introduce and analyze an efficient way of approximate sampling of distributions on the sphere.
Let $\mathbb{S}^{d-1}$ be the (d1)(d-1)-dimensional Euclidean unit sphere in $\mathbb{R}^{d}$ and let $\sigma_{d-1}$ denote the volume measure on $\mathbb{S}^{d-1}$. For $p\colon\mathbb{S}^{d-1}\to(0,\infty)$ satisfying
@@
    Posteriors naturally defined on the sphere require efficient sampling: To provide some examples, sampling distributions on the sphere play an important role in directional statistics and shape analysis (see Mardia and Jupp, 2000), and Bayesian inverse problems on $\mathbb{S}^{2}$ occur, e.g., in astrophysics or geophysics (see Marignier et al., 2023). Moreover, Bayesian density estimation (Holbrook et al., 2020) requires sampling of a posterior on an infinite-dimensional sphere, which usually is approximated by truncating the dimension, ending up with $\mathbb{S}^{d-1}$ for large $d$. Also the problem of 3D rigid registration can be reduced to sampling from a distribution on $\mathbb{S}^{3}$ (see Section 4.1).
@@
    Spherical sampling can be used as tool within transforming Markov chains: Lan et al. (2014) introduced a Hamiltonian Monte Carlo (HMC) scheme for sampling from spherical distributions that leads, after suitable transformations, to an efficient exploration of distributions in $\mathbb{R}^{d}$ constrained by inequalities. Moreover, recently Yang et al. (2024) demonstrated that MCMC algorithms on $\mathbb{S}^{d-1}$ can target distributions in $\mathbb{R}^{d}$ by mapping them to the sphere by means of the stereographic projection and then sampling from the transformed target on the sphere. This is further developed by Bell et al. (2025) with an adaptive MCMC strategy. The authors report that the “stereographic scheme” performs well in the presence of heavy-tailed distributions, as shown by both theoretical and empirical evidence. Bell et al. (2025) also indicate that the combination with slice sampling on the sphere, proposed here, is particularly effective and outperforms competing methods.
The advantage of the slice sampling paradigm is that it enables an automatic, location-dependent traversal of the space via density based regions, without requiring any laborious manual tuning of parameters to ensure efficient exploration of the state space. This usually results in an algorithm that is less sensitive to certain degenerations of the target distribution such as multimodality or anisotropy, see also Neal (2003); Murray et al. (2010). A transition mechanism for realizing a corresponding Markov chain works as follows: Given current state $x\in\mathbb{S}^{d-1}$, a superlevel set $L(t)$ of $p$ is randomly determined by choosing a level $t\in(0,p(x))$. Then, the next Markov chain instance is specified by (mimicking) sampling of the normalized reference measure $\sigma_{d-1}$ restricted to the level set. The latter step requires some care regarding its algorithmic design. The main contributions of our paper are:
  • We introduce an ideal geodesic slice sampler with corresponding Markov kernel $H$ for approximately simulating $\pi$. The kernel $H$ is implemented by moving along a randomly chosen great circle through the current state. The resulting sampling on the great circle (understood as geodesic) intersected with the level set can be algorithmically realized by using an univariate acceptance/rejection approach.
@@
  • We test our algorithms on challenging targets such as a posterior distribution arising in rigid registration and mixtures of von Mises-Fisher distributions. We observe that, in a variety of settings, our slice samplers outperform standard samplers such as random walk Metropolis Hastings (RWMH) and Hamiltonian Monte Carlo (HMC) on spherical distributions.
Let us comment on how our contributions fit into the literature. Lie et al. (2023) present two MCMC algorithms to sample distributions on the sphere using push forward kernels of a preconditioned Crank-Nicolson algorithm and of elliptical slice sampling, respectively. The crucial difference to our work is that Lie et al. (2023) assume that the target density is defined relative to an angular Gaussian distribution. Instead, here we consider target densities relative to the volume measure of the sphere. Observe that with increasing dimension both reference measures become “increasingly singular” to each other. Furthermore, Marignier et al. (2023) propose a proximal MCMC method to sample from posterior distributions of inverse imaging problems on $\mathbb{S}^{2}$. An infinite-dimensional setting of Bayesian density estimation is treated by Holbrook et al. (2020) by an HMC algorithm on a sphere. Other MCMC approaches have been developed on more general manifolds that usually cover $\mathbb{S}^{d-1}$ as a special case. This includes the HMC algorithm for manifolds embedded in the Euclidean space introduced by Byrne and Girolami (2013) which uses the geodesic flow, and the HMC algorithm for manifolds that are the fibre of a smooth map proposed by Brubaker et al. (2012). Furthermore, Goyal and Shetty (2019) investigate a Metropolis-Hastings-like geodesic walk on manifolds with non-negative curvature. Similarly, the Metropolis-Hastings algorithm of Zappa et al. (2018) based on projections from the tangent spaces can be applied to $\mathbb{S}^{d-1}$. Here it is also worth to mention that several geodesic walks exist that target the uniform distribution on subsets of the sphere. We refer to Section 2.3 for more details. Finally, note that recently there has been some theoretical progress in the investigation of convergence properties of slice sampling (see for example Łatuszyński and Rudolf, 2024; Natarovskii et al., 2021b, a; Hasenpflug et al., 2024) that very much influenced the presentation and proof arguments of our theoretical results.
@@ 2 Preliminaries
In this section we provide some general notation, give a brief introduction to geometrical features of the Euclidean unit sphere and explain how these can be leveraged to “walk the sphere”.
@@ 2.1 Setting and Notation
Let $\lambda$ be the Lebesgue measure on $\mathbb{R}$, and let $\left\|\cdot\right\|$ be the Euclidean norm induced by the standard inner product ${x}^{T}\!{y}$, where $x,y\in\mathbb{R}^{d}$ for $d\in\mathbb{N}$. Moreover, define $\mathrm{Id}_{d}\in\mathbb{R}^{d\times d}$ to be the $d$-dimensional identity matrix. Throughout this paper we assume $d\geq 3$ and consider the Euclidean unit sphere
@@
On a generic Borel space $(\mathcal{X},\mathcal{B}(\mathcal{X}))$ define the restriction of a measure $\nu$ to a set $A\in\mathcal{B}(\mathcal{X})$ by $\nu|_{A}$, that is $\nu|_{A}(B)=\nu(A\cap B)$ for all $B\in\mathcal{B}(\mathcal{X})$. Whenever appropriate, for a set $A\in\mathcal{B}(\mathcal{X})$ we denote by $\mathcal{U}_{A}$ the uniform distribution on $A$. For example, for $\mathcal{X}=\mathbb{R}$ with $A\in\mathcal{B}(\mathbb{R})$ the probability measure $\mathcal{U}_{A}$ refers to $\frac{1}{\lambda(A)}\,\lambda|_{A}$ if $\lambda(A)\in(0,\infty)$ and for $\mathcal{X}=\mathbb{S}^{d-1}$ the distribution $\mathcal{U}_{\mathbb{S}^{d-1}}$ refers to $\frac{1}{\omega_{d-1}}\sigma_{d-1}$, where $\omega_{d-1}\coloneqq\sigma_{d-1}\left(\mathbb{S}^{d-1}\right)$. Sometimes it is more convenient to work with random variables. Therefore, let $(\Omega,\mathcal{F},\mathbb{P})$ be a sufficiently rich probability space on which all random variables occurring in this paper are defined. If a random variable $X$ has distribution $\nu$, we write $X\sim\nu$.
@@ 2.2 The Geometry of the Sphere
the great subsphere with pole $x$. This (d2)(d-2)-dimensional subsphere is the intersection of $\mathbb{S}^{d-1}$ and the (d1)(d-1)-dimensional hyperplane perpendicular to $x$, and in this sense can be thought of as a generalization of the equator. We also equip $\mathbb{S}^{d-2}_{x}$} with the corresponding volume measure, which we denote as $\mu_{x}$. Since $\mathbb{S}^{d-2}_{x}$} is essentially “a tilted $\mathbb{S}^{d-2}$”, we have $\mu_{x}({\mathbb{S}^{d-2}_{x}})=\omega_{d-2}$ for all $x\in\mathbb{S}^{d-1}$, cf. Appendix A. Accordingly $\frac{1}{\omega_{d-2}}\mu_{x}$ coincides with $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}$. In our transition mechanism, sampling w.r.t. the latter distribution frequently appears, such that we provide a procedure for performing it in Appendix A.
@@
Intuitively, we can also think of $\gamma_{(x,v)}$ as the curve obtained by “moving from the point $x$ in the direction of $v$ with a constant velocity”. This interpretation originates from Riemannian geometry and following its terminology we may use the terms great circle and geodesic interchangeably for the object $\gamma_{(x,v)}$. Note that all great circles passing through a point $x\in\mathbb{S}^{d-1}$ are of the from $\gamma_{(x,v)}$ for some $v\in{\mathbb{S}^{d-2}_{x}}$, i.e., the great circles through $x\in\mathbb{S}^{d-1}$ are parametrized by $\mathbb{S}^{d-2}_{x}$}.
Observe that, due to the periodicity of sine and cosine, all great circles are $2\uppi$-periodic. This is exploited when we incorporate great circles into our slice sampling approach.
@@
for $\theta\in\mathbb{R}$, where for brevity we define $\mathcal{S}\mathbb{S}^{d-1}\coloneqq\bigcup_{x\in\mathbb{S}^{d-1}}\big{(}\{x\}\times{\mathbb{S}^{d-2}_{x}}\big{)}$. If we interpret for a pair $(x,v)\in\mathcal{S}\mathbb{S}^{d-1}$ the first component as position on the sphere and the second component as “direction of view”, then applying the map $T_{\theta}$ to $(x,v)$ can be thought of as “following the great circle $\gamma_{(x,v)}$ from $x$ in direction $v$ for a length of $\theta$ and then doing a U-turn back on the spot”. This intuition is underpinned by the following lemma, proven in Appendix B, which essentially tells us that following the great circle corresponding to the pair $T_{\theta}(x,v)$ is the same as following the original great circle $\gamma_{(x,v)}$ in reverse direction with an offset of $\theta$.
@@
For the first step in the outlined transition we can use the fact that great circles through $x\in\mathbb{S}^{d-1}$ are parametrized by $\mathbb{S}^{d-2}_{x}$}. Namely, we realize “a random great circle” as $\gamma_{(x,v)}$, where $v$ is a sample from $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}$, see Algorithm 5 in Appendix A for the simulation of this probability measure. The distribution of the randomly chosen point on the great circle in step 2 significantly influences the stationary distribution of the random walk on the sphere. In this section, we consider it to be independent of $x$, which causes the stationary distribution to be the uniform one. In Section 3, we explain how the second step can be modified so as to obtain a Markov chain with a desired stationary probability measure.
Using a distribution $\tau$ on $0,2\uppi)$ to sample on the great circle, we obtain the following Markov kernel
@@
1: Draw $V\sim\mathcal{U}_{\mathbb{S}^{d-2}_{x}}$, call the result $v$.   {Perform Algorithm 5 with input xx.}
@@
For targeting the uniform distribution on the sphere, the geodesic random walk introduced by Goyal and Shetty (2019, Algorithm 1, for their $K=\mathbb{S}^{d-1}$) corresponds to $K_{\tau}$ with $\tau$ being the distribution of $\varepsilon R$, where $R$ is a chi-distributed random variable with d1d-1 degrees of freedom and fixed $\varepsilon>0$. For manifolds with non-negative sectional curvature and bounded Riemannian curvature tensor, Goyal and Shetty (2019) provide mixing time results if their unfiltered walk is modified to target the uniform distribution of a strongly geodesically convex subset of the ambient manifold.
@@ 3 Geodesic Slice Sampling in $\mathbb{S}^{d-1}$
Proof  By the lower semicontinuity $p^{-1}((t,\infty))=L(t)$ is open, and by the fact that $x\in L(t)$, it is also non-empty. Moreover, since $\gamma_{(x,v)}$ is continuous, we have that $\gamma_{(x,v)}^{-1}(L(t))$ is again open and non-empty. Together with the $2\uppi$-periodicity of $\gamma_{(x,v)}$, this implies that $L(x,v,t)=\gamma_{(x,v)}^{-1}(L(t))\cap[0,2\uppi)$ has non-empty interior, such that $\lambda(L(x,v,t))>0$. The final assertion follows by observing that $L(\left\|p\right\|_{\infty})$ is open (by the lower semicontinuity) and by the definition of $\left\|p\right\|_{\infty}$ satisfies $\sigma_{d-1}(L(\left\|p\right\|_{\infty}))=0$. Consequently, we have $L(\left\|p\right\|_{\infty})=\emptyset$, such that $\left\|p\right\|_{\infty}=\sup_{x\in\mathbb{S}^{d-1}}p(x)$.